# CLEAR: Character Unlearning in Textual and Visual Modalities

### v1

@@ Abstract
Machine Unlearning (MU) is critical for enhancing privacy and security in deep learning models, particularly in large multimodal language models (MLLMs), by removing specific private or hazardous information. While MU has made significant progress in textual and visual modalities, multimodal unlearning (MMU) remains significantly underexplored, partially due to the absence of a suitable open-source benchmark. To address this, we introduce CLEAR, a new benchmark designed to evaluate MMU methods. CLEAR contains 200 fictitious individuals and 3,700 images linked with corresponding question-answer pairs, enabling a thorough evaluation across modalities. We assess 10 MU methods, adapting them for MMU, and highlight new challenges specific to multimodal forgetting. We also demonstrate that simple $\ell_{1}$ regularization on LoRA weights significantly mitigates catastrophic forgetting, preserving model performance on retained data. The dataset is available at https://huggingface.co/datasets/therem/CLEAR.
@@
Alexey Dontsov1,6, Dmitrii Korzh1,3, Alexey Zhavoronkin2,4, Boris Mikheev3, Denis Bobkov1,6, Aibek Alanov1,6, Oleg Y. Rogov1,3,5, Ivan Oseledets1,3, Elena Tutubalina1,6 1AIRI 2MIPT 3Skoltech 4Sber 5University of Sharjah 6HSE University Correspondence: dontsov@airi.net; tutubalina@airi.net
@@ 1 Introduction
Figure 2: Summary of our dataset. We generate 200 persons and use multimodal unlearning to forget the part of them. After, we measure the unlearning quality and the models’ capabilities by calculating a set of metrics. Then, we create a leaderboard of unlearning methods based on these metrics.
Large Language Models (LLMs) Touvron et al. (2023); Jiang et al. (2023) are trained on vast corpora of data that contain private, unethical, or unwanted information, leading to growing concerns. Machine unlearning (MU) methods have been developed to remove such unwanted data without expensive retraining from scratch. For instance, MU has been applied for the LLMs to mitigate issues related to toxicity Lu et al. (2022), copyright and privacy concerns Jang et al. (2022); Eldan and Russinovich (2023); Wu et al. (2023) and fairness Yu et al. (2023). Additionally, such topics as model editing Ilharco et al. (2022); Zhang et al. (2023), prevention of hallucinations Yao et al. (2023), and sensitive knowledge exposure Barrett et al. (2023) have also motivated the development of MU techniques.
There are various unlearning techniques suitable solely for LLMs Yao et al. (2024b, a); Xing et al. (2024); Zhang et al. (2024) or for vision models Li et al. (2024a); Chen and Yang (2023); Tarun et al. (2021). However, multimodal LLMs (MLLMs) Liu et al. (2023), specifically visual LLMs (VLLMs), raise new challenges. The unlearning of such multimodal models (MMU) remains largely unexplored, primarily due to the lack of open-source benchmarks. Moreover, current MU benchmarks Maini et al. (2024); Shi et al. (2024); Yao et al. (2024a); Li et al. (2024b) are focused on single modalities, and, to our knowledge, no open benchmarks designed explicitly for evaluating unlearning in multimodal models exist at the time of submission.
To address this gap, we propose CLEAR, a new benchmark for textual-visual MMU named, focusing on person unlearning, which aligns with the right-to-be-forgotten concept. The dataset is synthetic to ensure control over the data the model learns, preventing object leakage during training. We generated consistent images through a comprehensive strategy and linked them to the corresponding author-related questions from the large-scale textual unlearning benchmark TOFU Maini et al. (2024). The proposed dataset contains 200 fictitious authors, 3,770 visual question-answer pairs, and 4,000 textual question-answer pairs, enabling a thorough evaluation of the single and multi-modal unlearning techniques. We also propose a benchmark to assess MU and MMU methods, evaluating 10 techniques, including the current state of the art.
We evaluate the unlearning methods in textual, visual, and multimodal setups. First, we finetune the model on our entire dataset. Next, we take a small predefined subset of 20 authors, referred to as the forget set, while the remaining data forms the retain set. We then apply the unlearning procedure, resulting in a new model that no longer "remembers" individuals from the forget set but retains knowledge of those in the retain set. To ensure the model’s capabilities are not compromised during unlearning, we assess its performance on real-world tasks, such as celebrity face recognition and general domain visual question-answering (VQA). Overall, we work with four sets: Forget, Retain, Real Faces, and Real World. Figure 1 illustrates an example from each set.
We evaluate existing unlearning methods separately in textual and vision modalities and then combine them within an MLLM. We create a leaderboard for each domain, highlighting that multimodal unlearning poses new challenges.
Finally, we demonstrate that applying $\ell_{1}$ regularization to the LoRA adapter during the unlearning process significantly improves performance, helping to prevent catastrophic forgetting of the retain set information. Overall, our contributions can be summarized as follows:
  • We propose a novel benchmark, CLEAR, for evaluating machine unlearning in multimodal (textual-visual) setups. To the best of our knowledge, this is the first publicly available MMU benchmark.
  • We comprehensively evaluate existing unlearning methods across separate and combined domains. We construct leaderboards for these three domains and show that state-of-the-art unlearning algorithms struggle in multimodal setups, highlighting the need for new approaches.
  • We demonstrate that the $\ell_{1}$ weight regularization on the LoRA adapter helps to improve unlearning quality by significantly preventing catastrophic forgetting.
2 Related Work
2.1 MU Methods and Textual Benchmarks
MU methods Cao and Yang (2015); Dwork et al. (2014); Kurmanji et al. (2024); Neel et al. (2021); Sekhari et al. (2021) remove the impact of certain data instances from a trained model without requiring full retraining. The goal is to obtain a model that behaves like the forget data was never part of the training set. MU can be formalized in two main ways. In a setting where the unlearned model must produce identical outputs to a model trained without the forget data. In an inexact unlearning setting, the main objective is to obtain any model that doesn’t contain knowledge from the forget set, but with no restrictions and guarantees on the model’s response to inputs from the retain set.
There are several standard textual unlearning benchmarks. TOFU Maini et al. (2024) is a benchmark for textual LLM unlearning, featuring 200 fictitious author profiles, each defined by attributes such as name, birthplace, parent’s names and occupation, written books, etc. Totally the dataset has 4,000 question-answer pairs (20 per author). Predefined 10/90, 5/95, and 1/99 splits are used for forget/retain sets. Models are finetuned on the entire dataset, and unlearning is applied to the forget pairs. WMDP Li et al. (2024b) consists of 3,668 multiple-choice questions, evaluates hazardous knowledge of LLMs, and serves as a benchmark for unlearning such dangerous information. In (Yao et al., 2024a), the authors explore seven textual unlearning techniques. The best approach was a combination of gradient ascent with gradient descent on the forget and retain sets, respectively, to maintain the unlearn/retain quality trade-off. They evaluated their methods on three benchmark datasets from different domains: 500 arXiv papers, 2000 GitHub files, and 100 books covering academic texts, code, and literary works. Unfortunately, these three benchmarks cannot be applied to the MMU evaluation.
2.2 MMU Methods and Benchmarks
MLLMs Liu et al. (2023) typically consist of three core components: a modality encoder that translates raw input into feature embeddings, a modality projection layer aligning these features within the language space, and a pre-trained language model synthesizing the final output.
Nonetheless, MMU research is still in its early stages. Cheng and Amiri (2023) proposed the MultiDelete method, which focuses on separating cross-modal embeddings for the forget set while preserving unimodal embeddings for the retain set. Unfortunately, this approach is suitable only for the encoder-decoder architecture and can not be directly transferred to decoder-only LLMs.
EFUF Xing et al. (2024) mitigates hallucinations in MLLMs using unlearning. It measures the similarity between generated captions and image content with CLIP model Radford et al. (2021) to automatically detect hallucinated (negative) and non-hallucinated (positive) examples based on the calibrated on the MSCOCO dataset Lin et al. (2014) thresholds, eliminating manual labeling. The unlearning process applies three loss functions: negative loss to forget hallucinations, positive loss to reinforce correct representations, and sentence loss to maintain fluency. However, their benchmark is not open-sourced.
Single Image Unlearning (SIU) Li et al. (2024a) focuses on unlearning visual concepts in MLLMs while preserving textual knowledge and introduces MMUBench with five evaluation metrics. The benchmark covers 20 concepts with at least 50 images for the concept, including real-world figures and cartoon characters. For each concept, one image is selected as the forget subset, paired with various prompts, while the remaining images form the retain subset. However, SIU’s use of a single image raises scalability concerns for complex concepts, and their unlearning of VLLM is limited to the visual domain. Moreover, MMUBench is also not open-sourced.
Chakraborty et al. (2024) explores unlearning harmful content in VLLMs, demonstrating that unlearning in the textual domain alone can match the performance of text-image unlearning while using fewer resources. Their method combines harmful loss and KL divergence between unlearned and retrained models. The approach was tested on six datasets, including PKU-SafeRLHF Ji et al. (2024) and three vision-text attack datasets Shayegani et al. (2023); Luo et al. (2024); Gong et al. (2023) for harmful content prompts, and Truthful-QA and VQA-v2 Lin et al. (2021); Goyal et al. (2017) to ensure benign task performance remained intact. However, their focus on safety alignment may limit applicability for general unlearning, such as biometric privacy. While they claim that textual unlearning is sufficient for MMU, our findings show this is not true for all methods. Additionally, their approach lacks exact unlearning evaluation.
3 Methodology
Unlearning can be conceptualized in two main ways: the objective of the Strict Unlearning is to achieve a model that behaves identically to one trained exclusively on the retain set, ensuring that no knowledge from the forget set is present; the objective of the Inexact Unlearning is to produce any model that no longer contains information from the forget set. However, this approach cannot guarantee how the model will respond to inputs related to the forget set. These objectives require distinct methods, loss functions, and evaluation metrics to assess unlearning quality.
Let $f_{\theta}$ denote the original model with its parameters $\theta$. Source model $f_{\theta}$ is trained on train dataset $D$, and given the unlearning objective, we want to make our model forget a subset of the source dataset $D$, called forget set $D_{F}$. The remaining part of the training dataset is called retain set, and we aim to preserve the model’s performance on this data subset $D_{R}:=D\setminus D_{F}$. Additionally, we utilize a holdout set $D_{H}$ to establish a reference for the model’s desired behavior on $D_{F}$ after the unlearning process. The model’s training did not include this set, ensuring that $D_{H}\cap D=\emptyset$. In a nutshell, forget set $D_{F}$ contains samples the model should unlearn and serves as a direct measure of unlearning effectiveness; retain set $D_{R}$ contains samples that the model should retain and perform well on, serving as an indicator of the model’s preserved knowledge; holdout set $D_{H}$ contains samples that the model has never seen before and serves as a reference for the model’s behavior on data that was not involved in the training process. Such forgetting procedure is performed by updating the model $f_{\theta}$ with a particular unlearning method, which results in a new unlearned model $f_{\hat{\theta}}$ with parameters $\hat{\theta}$. For the evaluation, we can also train separately a gold model $g_{\omega},$ which is trained only on the $D_{R}$.
So the objective is to obtain the unlearned model $f_{\hat{\theta}}$ which unlearns the forget set $D_{F}$ while preserving the performance on retain set $D_{R}$ as good as the source model $f_{\theta}$. It can be done by optimizing (minimizing or maximizing) the specific criterion with the model parameters $\theta$ and resulting obtained parameters $\hat{\theta}$ will be the parameters of desired unlearned model $f_{\hat{\theta}}$. For example, one can consider the gradient difference MU approach for the optimization, aimed at increasing forget loss and maintaining retain performance:
$\tilde{L}=-\sum_{x_{i}\in D_{f}}L(x_{i},y_{i},\theta)+\lambda\sum_{x_{j}\in D_{R}}L(x_{j},y_{j},\theta)$ (1)
$\theta\mapsto\theta-\alpha\nabla_{\theta}\tilde{L},$ (2)
where $\lambda$ – forget-retain trade-off hyper-parameter, $\alpha$ – learning rate, $L$ is a loss function, for example, negative-log-likelihood. $x$ is the input – text, image, or both of them in the case of VLLM.
In this work, we explore a subset of unlearning methods, including Retain Finetune, Gradient Ascend (GA), Gradient Difference (GD) Liu et al. (2022), SCRUB Kurmanji et al. (2023), DPO Rafailov et al. (2023), NPO Zhang et al. (2024), LLMU Yao et al. (2024b) , IDK Maini et al. (2024), RMU Li et al. (2024b), and KL Golatkar et al. (2020), which are detailed in Appendix A.We selected these methods based on their ease of adaptation to new modalities, requiring only changes in input data (text, images, or both) while preserving their core functionality. In essence, these methods involve variations of hard negative training on the forget set combined with fine-tuning on the retain set, often with additional constraints such as Cross-Entropy or KL divergence to align the model’s outputs on the retain set with the original model.
4 CLEAR
The MU (and consequently MMU) benchmark should ideally avoid running unlearning on well-known information that could be obtained from external sources such as books, games, movies, etc. This is essential for a more reliable evaluation of the model’s performance on retain and forget. To meet this requirement, we chose to extend the TOFU dataset due to its ease of use, flexibility for adopting to new modalities (such as adding face images or personal voices), and its strong connection to privacy concerns, making it ideal for testing unlearning in sensitive contexts.
4.1 Dataset Generation Process
Firstly, for each of the 200 authors from the TOFU dataset, we extract their name, age, and ethnicity based on the knowledge provided in the original dataset. Also, we generate a pool of 2000 faces using StyleGAN2 Karras et al. (2020) - an established generative model for face synthesis. Each face is scored with pre-trained CNNs to get the age, gender, and ethnicity of the face. Then, we manually select the most suitable face that matches these attributes for each author. During this phase, we discovered that the age distribution of the authors was highly shifted towards the older ages, so we needed to eliminate this gap. To do this, we used the image editing framework proposed in Bobkov et al. (2024) to shift the visual attributes of the faces to make them older. The distribution of the characteristics of the faces and authors is shown in Figure 4. After matching each author to a face, we used the diffusion model Li et al. (2024c) of personalized generation to synthesize images with a given face and corresponding to a given prompt. In detail, the face generation and collection process is described in Appendix B.
@@ 4.1 Dataset Generation Process
The diffusion model needs a textual prompt for image generation besides the face. We ask GPT-4 to generate images from a textual question and an answer about an author. We generate 8 images for each prompt, evaluate them using an ensemble of fake-detection models, and select the most realistic. Additionally, GPT-4o generates captions for each image and visual prompt pair, which are then included in the dataset. However, due to restrictions caused by GPT guard breaks and the identification of several bugs in the TOFU dataset (such as a nameless author), the final dataset includes fewer images than text pairs (3,770 compared to 4,000). We also incorporated two additional data splits containing real-world face photos and natural images to maintain the model’s visual capabilities during unlearning.

### v3

@@ Abstract
Machine Unlearning (MU) is critical for removing private or hazardous information from deep learning models. While MU has advanced significantly in unimodal (text or vision) settings, multimodal unlearning (MMU) remains underexplored due to the lack of open benchmarks for evaluating cross-modal data removal. To address this gap, we introduce CLEAR, the first open-source benchmark designed specifically for MMU. CLEAR contains 200 fictitious individuals and 3,700 images linked with corresponding question-answer pairs, enabling a thorough evaluation across modalities. We conduct a comprehensive analysis of 11 MU methods (e.g., SCRUB, gradient ascent, DPO) across four evaluation sets, demonstrating that jointly unlearning both modalities outperforms single-modality approaches.
@@
Alexey Dontsov1,6, Dmitrii Korzh1,3, Alexey Zhavoronkin2,4, Boris Mikheev3, Denis Bobkov1,6, Aibek Alanov1,6, Oleg Y. Rogov1,3,5, Ivan Oseledets1,3, Elena Tutubalina1,6 1AIRI 2MIPT 3Skoltech 4Sber 5MTUCI 6HSE University Correspondence: dontsov@airi.net; tutubalina@airi.net
@@ 1 Introduction
Large Language Models (LLMs) Ouyang et al. (2022); Touvron et al. (2023); Jiang et al. (2023) are increasingly investigated for memorizing private, unethical, or copyrighted data during training. Recently, machine unlearning (MU) methods have been applied to mitigate issues related to toxicity Lu et al. (2022), copyright and privacy concerns Jang et al. (2023); Eldan and Russinovich (2023); Wu et al. (2023) and fairness Yu et al. (2023).
Figure 1: Our dataset CLEAR includes 3,770 visual image-caption pairs and 4,000 textual question-answer pairs related to fictional characters. We apply multimodal unlearning to remove specific information, subsequently assessing the quality of unlearning and the models’ performance using various metrics. Finally, we compile a leaderboard of unlearning methods based on these evaluations.
While MU has emerged as a promising alternative to costly retraining—enabling targeted removal of problematic data—existing methods focus almost exclusively on single-modality models. Recent work has studied unlearning in LLMs Yao et al. (2024c, a); Xing et al. (2024); Zhang et al. (2024a) or vision models Li et al. (2024a); Chen and Yang (2023); Tarun et al. (2021), but unlearning in multi-modal language models remains largely unexplored. This leaves a critical gap: multimodal LLMs (MLLMs), which process both visual and textual data, introduce unique challenges for unlearning. For instance, sensitive information may persist across modalities even after removal from one (e.g., a face linked to a name), and unlearning in one modality could degrade performance in another. Despite these risks, no open benchmarks exist to evaluate MU in multimodal settings.
Method Real $\uparrow$ Retain metric $\uparrow$ Forget metric $\downarrow$ Forget Quality $\uparrow$
Gold 0.50 0.51 0.19 1.00
Base 0.48 0.51 0.35 0.85
DPO 0.46 0.48 0.22 0.84
GD 0.29 0.00 0.00 0.18
GA 0.27 0.00 0.00 0.67
IDK 0.48 0.51 0.33 0.84
KL 0.25 0.00 0.00 0.67
LLMU 0.47 0.51 0.25 0.84
NPO 0.46 0.14 0.11 0.76
Retain FT 0.49 0.51 0.37 0.85
RMU 0.24 0.00 0.00 0.75
SCRUB 0.49 0.52 0.36 0.85
SKU 0.40 0.32 0.37 0.83
Table 1: Performance comparison of state-of-the-art unlearning methods on our dataset across four metrics. “Base” refers to the model before unlearning, while “Gold” denotes a model trained only on the retain set. The highlighted methods fail on the retain set.
Recently, Chakraborty et al. (2024) pioneers the investigation of unlearning configurations in visual-language models (VLMs) to mitigate cross-modal safety risks, its experimental framework inherits a critical limitation: the datasets used (e.g., PKU-SafeRLHF Ji et al. (2023), JailBreakV-28K Luo et al. (2024)) were designed for safety alignment and truthfulness evaluation, not machine unlearning (MU). This mismatch conflates safety fine-tuning (suppressing harmful outputs) with targeted data removal (erasing specific knowledge traces), potentially overestimating MU efficacy.
To address this, we propose CLEAR, the first publicly available benchmark for machine unlearning in multimodal (textual-visual) models. Our work is motivated by the right to be forgotten in AI systems, where models must eliminate traces of specific entities (e.g., individuals) across all modalities. The proposed dataset contains information about fictitious authors. Each persona is linked to both textual biographies and AI-generated images, enabling tests of cross-modal memorization. After unlearning a persona, models should fail to answer questions and recognize associated faces. Our benchmark further evaluates real-world performance degradation using visual question-answering (VQA) tasks.
Our contributions and findings are as follows:
  • We propose a multimodal MU benchmark CLEAR with 4,000 text-QA pairs and 3,770 image-caption pairs focused on unlearning 200 fictitious authors. It includes forget/retain sets and real-world tasks (e.g., celebrity recognition) to evaluate cross-modal capability preservation.
  • We comprehensively evaluate 11 recently proposed MU methods on our dataset and show that leading unimodal MU methods struggle in multimodal setups.
  • We establish leaderboards for textual, visual, and multimodal unlearning.
We make all the data publicly available111at https://huggingface.co/datasets/fd28/CLEAR.
2 Related Work
Machine Unlearning. The concept of machine unlearning was initially presented by Cao and Yang (2015). In general, MU methods Cao and Yang (2015); Dwork et al. (2014); Kurmanji et al. (2024); Neel et al. (2021); Sekhari et al. (2021) remove the impact of specific data points from a trained model without requiring full retraining. The goal is to obtain a model that behaves like the forget data was never part of the training set. Recently, textual unlearning in generative language models has attracted attention. Maini et al. (2024) propose a benchmark named TOFU for textual LLM unlearning, consisting of 200 fictitious author profiles defined by attributes like name, birthplace, parent’s names, occupation, and written books, totalling 4,000 question-answer pairs (20 per author). WMDP Li et al. (2024b) includes 3,668 multiple-choice questions to evaluate and benchmark the unlearning of hazardous knowledge in LLMs. To remove knowledge from generative models, Jang et al. (2023) employ gradient ascent on specific target sequences. Eldan and Russinovich (2023) focus on the particular case of unlearning the Harry Potter books from Llama2-7b. Yao et al. (2023) utilize machine unlearning to address harmful responses and eliminate hallucinations. Yao et al. (2024b) examined the unlearning of 2,000 GitHub code files, 500 books, and 500 academic papers from Yi-6B. However, these studies have been restricted to text-only contexts. Our research investigates the multimodal aspects of unlearning.
Multimodality.
Multimodal LLMs Liu et al. (2023) usually comprise a modality encoder, a projection layer aligning features to the language space, and a pre-trained language model. While MLLMs have advanced, multimodal unlearning remains under-explored. Cheng and Amiri (2023) introduce MultiDelete, which separates cross-modal embeddings for unlearning but applies only to encoder-decoder models, limiting its use for decoder-only architectures. EFUF Xing et al. (2024) reduces hallucinations in MLLMs by unlearning. It uses CLIP Radford et al. (2021) to detect hallucinations based on MSCOCO-calibrated thresholds, eliminating manual labelling. The method applies three losses: negative loss to forget hallucinations, positive loss to reinforce correct representations, and sentence loss to preserve fluency. Single Image Unlearning (SIU) Li et al. (2024a) targets visual concept unlearning in VLLMs while preserving textual knowledge and introduces MMUBench. This benchmark spans 20 concepts with 50+ images each, including real-world figures and cartoon characters. However, these benchmarks are not open-sourced. The closest work to ours is Chakraborty et al. (2024), which investigates safety alignment in VLMs by unlearning harmful content. This study was not designed for the exact unlearning setup; therefore, we bridge this gap by conducting the first comprehensive analysis of MU methods in multimodal settings. Unlike safety alignment, our benchmark focuses on forgetting quality (e.g., inability to recall personas) and cross-modal consistency (e.g., erasing both a face and its biography) while maintaining model utility in real-world tasks, highlighting unique challenges in multimodal MU.
3 MU Methods
3.1 Preliminaries
Let $f_{\theta}$ denote the base model with parameters $\theta$. It is trained on (train) dataset $D$, and given the unlearning objective, we want to make our model forget a subset of this dataset $D$, called forget set $D_{F}$. The remaining part of the training dataset is called retain set, and we aim to preserve the model’s performance on this data subset $D_{R}:=D\setminus D_{F}$. Additionally, we utilize a holdout set $D_{H}$ such that $D_{H}\cap D=\emptyset$ to establish a reference for the model’s desired behaviour on $D_{F}$ after the unlearning process. In a nutshell, forget set $D_{F}$ contains samples the model should unlearn and serves as a direct measure of unlearning effectiveness; retain set $D_{R}$ contains samples that the model should retain and perform well on, serving as an indicator of the model’s preserved knowledge; holdout set $D_{H}$ contains samples that the model has never seen before and serves as a reference for the model’s behaviour on data that was not involved in the training process. Such forgetting procedure is performed by updating the model $f_{\theta}$ with a particular unlearning method, which results in a new unlearned model $f_{\hat{\theta}}$ with parameters $\hat{\theta}$. Evaluation of $f_{\hat{\theta}}$ on the discussed subsets (or particularly on forget set) is called ”inexact” in contrast to the ”exact” evaluation when we directly compare the performance of the unlearned model with a gold model $g_{\omega}$, trained only on the $D_{R}$.
MU can be performed by optimizing the specific criterion. For example, one can consider the gradient difference MU approach, aimed at increasing forget loss and maintaining retain performance:
$\tilde{L}=-\sum_{x_{i}\in D_{f}}L(x_{i},y_{i},\theta)+\lambda\sum_{x_{j}\in D_{R}}L(x_{j},y_{j},\theta),$ (1)
$\theta\mapsto\theta-\alpha\nabla_{\theta}\tilde{L},$ (2)
where $\lambda$ – forget-retain trade-off hyper-parameter, $\alpha$ – learning rate, $L$ is some loss function, e.g., negative-log-likelihood, $x$ is an input (text, image, or both of them in the case of VLLM).
Suppose that we are given the LLM model denoted as $f$ described by its parameters $\theta$, hence representing a function mapping the input to the corresponding prediction, as described below:
$f_{\theta}(x)=\prod_{i=1}^{|y|}P_{\theta}\left(y_{i}\mid y_{<i},x\right),$ (3)
where $P_{\theta}$ is the probability function for generating the next token in the given sequence $=(y_{1},\ldots,y_{\lvert y\rvert})$, and $y_{<i}=\{y_{1},\dots,y_{i-1}\}$. Given an unlearned descriptor $\left(x_{u},y_{u}\right)$ related to an unlearning instance $\mathcal{I}$ (e.g., public figures or copyright-protected information). Current approaches often indiscriminately update $\theta$ to $\theta^{\prime}$ to ensure that all responses, $y_{u}^{\prime}=f_{\theta^{\prime}}\left(x_{u}\right)$, related to $\mathcal{I}$ are non-harmful. Yet, not all knowledge tied to $\mathcal{I}$ necessarily is required to be forgotten in this process.
In multi-modal unlearning, compared to unimodal unlearning, each sample contains multiple modalities, e.g., text and image pairs. Let:
$D\;=\;\bigl{\{}\,\bigl{(}x_{i}^{\text{(1)}},\dots,x_{i}^{\text{(m)}},y_{i}\bigr{)}\bigr{\}}_{i=1}^{N},$ (4)
where $m\geq 2$ is the number of modalities. We similarly define the forget set $D_{F}$ of multi-modal samples that we aim to unlearn. The model $f_{\theta}$ is now capable of taking multi-modal input (e.g., a textual prompt plus an image) and producing an output $y$. The same forget and retain goals hold.
3.2 Methods
We briefly describe 5 top-performing MU methods from Tab. 1 among all approaches described in detail in Appx. B.
Retain Finetune
is a straightforward approach which involves finetuning the model on the retain set, assuming it will forget the knowledge from the $D_{F}$ while maintaining performance on the $D_{R}$. However, it is suboptimal for models with extensive pretraining, such as most LLMs.
IDK tuning
replaces original labels in the forget set $D_{F}$ with ”I don’t know” responses while minimizing the loss on the retain set $D_{R}$ Maini et al. (2024). The objective, $L_{\text{IDK}}=L(D_{R},\theta)+L(D_{F}^{\text{idk}},\theta)$, ensures that the model retains performance on $D_{R}$ while aligning predictions on $D_{F}$ with uncertainty-based responses.
LLMU
was introduced in early LLM unlearning research Yao et al. (2024c), optimizes the loss:
$L_{\text{LLMU}}=-L(D_{F},\theta)+L(D_{F}^{\text{idk}},\theta)\\ +\sum_{x,y\in D_{R}}\operatorname{KL}(p_{\theta}(y|x)||p_{\hat{\theta}}(y|x)).$
Here, $\theta$ and $\hat{\theta}$ are model’s parameters before and after unlearning, the first term promotes unlearning by maximizing loss on $D_{F}$, while the second reinforces forgetting using ”I don’t know” labels instead of original targets. The KL-divergence term preserves performance on the retain set $D_{R}$ by aligning model outputs before and after unlearning.
SCRUB (Teacher-Student)
formulates unlearning as a teacher-student setup, where a student model learns from a fixed teacher Kurmanji et al. (2023). The student is optimized to match the teacher on the retain set $D_{R}$ while deviating on the forget set $D_{F}$. The loss function combines KL-divergence for retention ($L_{R}$), enforced divergence for unlearning ($L_{F}$), and task loss ($L_{\text{task}}$), yielding $L_{\text{SCRUB}}=L_{R}-L_{F}+L_{\text{task}}$.
Preference Optimization (DPO)
method applies Direct Preference Optimization (DPO) Rafailov et al. (2023) for MU and MMMU. The model is trained to reduce reliance on undesired information through a loss function combining task performance retention ($L_{\text{task}}$) and a DPO-based loss ($L_{\text{DPO}}$), which penalizes deviations from a reference model $\pi_{ref}$ fine-tuned on $D_{F}^{\text{idk}}$ with ”I don’t know” labels as in IDK-tuning.
To sum up, we chose recently published MU methods for their easy adaptation to new modalities, needing only input data changes (text, images, or both) while maintaining core functionality.
4 CLEAR
In this section, we describe a new benchmark CLEAR designed for character unlearning. As a basis, our dataset utilizes the text-only TOFU dataset Maini et al. (2024) within the same experimental setup to replicate a real-world scenario where privacy concerns arise in sensitive contexts. While external information from books, games, or movies is general knowledge to unlearn Eldan and Russinovich (2023); Li et al. (2024a); Xing et al. (2024), character unlearning deals with specific contextual data that directly impacts individuals. This task addresses removing personal or confidential details and enhancing user privacy.
4.1 Dataset Generation Process
The generation of synthetic faces for author profiles in our benchmark is motivated by ethical, technical, and practical considerations (see the complete rationale in Appx. A). Firstly, for each of the 200 authors from the TOFU dataset, we extract their name, age, and ethnicity based on the knowledge provided in the original dataset. Also, we generate a pool of 2000 faces using StyleGAN2 Karras et al. (2020) - an established generative model for face synthesis. Each face is scored with a pre-trained image model to determine age, gender, and ethnicity. Then, for each author, we filter a pool of faces with similar characteristics and select the most appropriate one. We found out from textual information that the age distribution of the authors was highly shifted towards the older age group, so we needed to eliminate the age gap between authors’ profiles and corresponding images. To do this, we used the image editing framework proposed in Bobkov et al. (2024) to shift the visual attributes of the faces to make them older. The final distribution of face and author characteristics is shown in Fig.  2. After matching each author with a face, we used the diffusion model Li et al. (2024c) to synthesize images based on the given face and corresponding prompt (Appx.  C).
@@ 4.1 Dataset Generation Process
We perform a simple reality check to ensure the quality of generated faces. We use the CLIP ViT-L/14 model, usually considered a visual encoder for most of the VLMs, to get embeddings of these three image sets – our faces, CelebA Liu et al. (2015) and WebFace Yi et al. (2014). Then, we calculate the pairwise FID scores on top of the embeddings of these sets, and we get the following results: FID between our faces and CelebA is 74.4, between our faces and WebFace is 69.2, and between CelebA and WebFace is 62.1. This shows that the distance between our faces and the real-world faces is comparable to the distance between two real-world face datasets. Besides the author’s face, the diffusion model needs a textual prompt to produce an image. We ask GPT-4 to generate these prompts from a question-answer pair from TOFU about an author. We generate 8 images for each prompt, evaluate them using an ensemble of fake-detection models, and select the most realistic one. Additionally, GPT-4o generates captions for each (image, visual prompt) pair, which are then included in the dataset to form pairs (image, caption). However, due to restrictions caused by GPT guard breaks and the identification of several bugs in the TOFU dataset (such as a nameless author), the final dataset includes fewer images than text pairs (3,770 compared to 4,000).